OpenStudy (anonymous):

HELLPPP... In a triangle, the second angle is twice the measure of the first, and the third angle is three times the measure of the second. Find the measure of the third angle. A. 2O DEGREES B. 30 DEGREES C. 120 DEGREES D. 180 DEGREES

OpenStudy (ranga):

Let first angle = x second angle = 2x third angle = 3 * 2x = 6x Three angles of a triangle add up to 180: x + 2x + 6x = 180 solve for x then substitute and find the third angle.

OpenStudy (anonymous):

im so confused can you explain more of how to solve it?

OpenStudy (anonymous):

anyoneeee

OpenStudy (anonymous):

In a triangle there are 180 degrees. Use what @ranga showed you. x + 2x + 6x = 180 9x = 180 x = 20 Now, you are looking for the THIRD angle, which is 6x. So you plug x into 6x, which then becomes 6 * (20), which is 120 degrees, your final answer.